OpenStudy (anonymous):

solve by completing the square: x^2+8x+6=0

OpenStudy (anonymous):

this is (x+4)^2-10=0 (x+4)^2=10 x+4=+-sqrt(10) x=+-sqrt(10)-4

OpenStudy (anonymous):

can you explain this one to me

OpenStudy (anonymous):

(x+4)^2 is the same as x^2+8x+16 if we substract 10, it will be the same as x^2+8x+6 so (x+4)^2-10 is the same as x^2+8x+6 we can write our equation in (x+4)^2-10=0 form and solve it

OpenStudy (anonymous):

okay that makes sense, what do you do after the last step you have up there

OpenStudy (anonymous):

10 goes to the other side of the equation (x+4)^2=10 and we do square root (x+4)=+-sqrt(10) so x=+sqrt(10)-4 or x=-sqrt(10)-4 that's the answer

OpenStudy (anonymous):

oh okay, you make this seem so simple haha, I dont know why im having so much trouble with it, thank you.

OpenStudy (anonymous):

you're welcome